# Neil A. Chriss

### v1

@@ 1 Introduction
In some circumstances trading is a strategic interaction between market participants buying or selling the same asset over the same period of time. This is the case when firms are concerned with the collective impact of their trades on market prices and choose how and when to trade in order to minimize their overall cost of trading taking into consideration what the other firms are likely to do. The general situation just described, henceforth called trading in competition, was studied in [Chr24b] and [Chr24a] and there it was shown that there are unique Nash equilibria (in the sense of [Nas50], [Nas51] and Definition 2.2 below) that only depend on the number of firms involved, their target quantities and the market impact model the firms use to select their trades.
A key assumption and in our view a major shortcoming of our previous work is that it implicitly assumes trading in competition is a game with complete information, one for which participants are in possession of full knowledge of all relevant aspects of the game, which in this context means the number of firms, their target quantities and the market impact model they use. By contrast, in most real-world situations, trading in competition is a game with incomplete information, in which some or all participants lack full knowledge of at least some aspects of the game. In the case of trading, firms do not know the precise target quantities their competitors seek to trade, and are far from sure of the parameters of the market impact models they use or even which models they use.
@@
For more on von Neumann and Morgenstern’s treatment of incomplete information, see [Mye04]. The main point for this discussion is that incomplete information is a persistent feature in trading situations due to institutional, regulatory, and competitive factors.
The present paper aims to demonstrate how uncertainty can be incorporated into competitive trading using the framework of games with "Bayesian" players (or Bayesian games), as developed in John Harsanyi’s seminal work on games with incomplete information [Har67, Har94]. Harsanyi’s contributions provide a general framework for addressing information uncertainty in game theory and extend Nash’s equilibrium concept to settings involving incomplete information.
Building on this foundation, we extend the work on trading in competition presented in [Chr24a] and [Chr24b] by allowing for incomplete information. To achieve this, we first review the theory of games with Bayesian players in Section 2, where we detail the methods Harsanyi introduced for handling incomplete information. We then apply these concepts to the context of competitive trading.
In Section 3.1, we revisit the fundamental game-theoretic concepts employed in [Chr24a] and [Chr24b], providing the necessary groundwork for our analysis. Section 2 introduces the framework of Harsanyi’s Bayesian games, setting the stage for an in-depth exploration of incomplete information within the trading context. A central result of this approach is a method for transforming a game with incomplete information into an equivalent game of complete information. This transformation leverages Harsanyi’s framework to ensure that the resulting game retains the essential features of the original but can be analyzed as a complete information game.
The advantage of this transformation is significant: complete information games are guaranteed to have at least one Nash equilibrium [^1] It is a fundamental result that all $n$-player non-cooperative complete information games have at least one Nash equilibrium [Nas51].. The equilibria derived in the transformed game can then be mapped back to the original incomplete information game, yielding strategies that are not only viable but also exhibit properties interpretable as optimal within the context of competitive trading.
@@
The main result of this paper is a description of $n$-firm equilibria in trading with incomplete information, building on the framework introduced in [Chr24a]. This framework is based on the following assumptions, which are reviewed in detail in Section 3.3:
@@
    There are $n$ firms trading in competition with one another over a fixed and common period of time (the "trading period").
@@
    Each firm uses a market impact model based on the Almgren-Chriss framework (see [AC01]), a linear model with both permanent and temporary impact. This model features a single free parameter for the market impact.
The novelty of this paper lies in relaxing a key implicit assumption from [Chr24a]: that each firm’s target quantity and market impact parameter are common knowledge. In game theory, a fact is considered common knowledge among players if (i) every player knows the fact, (ii) every player knows that every other player knows the fact, (iii) every player knows that every other player knows that every player knows the fact, and so on. The concept of common knowledge was explicitly introduced to game theory only in the 1970s [^2] The term ”common knowledge” may have originated in David Lewis’ philosophical treatise [Lew08] and was formally defined in game theory by Aumann in his seminal work [Aum76], though it can be argued that the idea was implicitly present in earlier works, such as those of von Neumann and Morgenstern..
@@ 2.3 Bayesian players and player types
Consider an $n$-player game where player $i=1,\dots,n$ has $K_{n}$ types. We have the following definitions.
@@
    Type profile: a type profile is a set of types $y=\{t_{1}^{k_{1}},\dots,t_{n}^{k_{n}}\}$, one for each player $i=1,\dots,n$. Write $t_{i}^{k}\in y$ if type $i$ for player $k$ is represented in $y$.
@@
    Profile set: write $\mathcal{T}$ for the set of all type profiles for this game.
@@
    Restricted profile set: fix a player $i$ and type $k$, then write $\mathcal{T}_{i}^{k}$ for the set of all type profiles $y_{i}^{k}\in\mathcal{T}$ such that $t_{i}^{k}\in y_{i}^{k}$.
@@ 2.4 Strategy profiles and payoff functions
We begin by defining the key analog of a strategy in a game with incomplete information, referred to a strategy profile.
@@
    Strategy profile for a player: For player $i=1,\dots,n$, let $s_{i}^{*}=(s_{i}^{1},\dots,s_{i}^{K_{i}})$ be a tuple of strategies, where $s_{i}^{k}$ corresponds to the strategy associated with type $t_{i}^{k}$ for player $i$.
@@
    Strategy profile for all types in the game: Given strategy profiles $s_{i}^{*}$ for each of the $n$ players, we define $s^{*}=(s_{1}^{*},\dots,s_{n}^{*})$ as the combined strategy profile for all players across all types.
@@
    Strategies: Each player is equipped with strategies $s_{i}^{*}$ consisting of a tuple $(s_{i}^{1},\dots,s_{i}^{K_{i}}$ such that $s_{i}^{k}$ is an admissible strategy for player $i$ when represented as type $k$, We assemble all of the strategy profiles into a single strategy profile $s^{*}$. Thus each player’s strategy in this game is a strategy profile $s_{i}^{*}$ consisting of admissible strategies in the original game;
@@ 2.5 Bayesian equilibrium for two firms
We now specialize type theory and Bayesian equilibrium to two firm, illustrating how the theory works in the simpler case. Let $i=1,2$ be the firms and $K$, $M$ be the number of types for players 1 and 2 respectively. For a given firm $i$, ${\sm i}$ will denote the "other player" so that if $i=1$, ${\sm i}=2$ and if $i=2$, ${\sm i}=1$. We write $\pi_{i}^{k}(m)$ for the probability that firm ${\sm i}$ is type $m$ conditional on firm $i$ being type $k$.
Strategy profiles for players as in definition 2.5 now simplifies in the case of two players to a $K$-tuple of admissible strategies for player 1 and an $M$-tuple for player 2 as follows.
@@
and a strategy profile for all types is given by a vector concatenating $s_{1}^{*}$ and $s_{2}^{*}$:
@@
Type profiles, Definition 2.4 are sets of types one for each player. There are $M\times N$ type profiles, and these can be written in lexicographical order as follows:
$y_{1}=(t_{1}^{1},t_{2}^{1}),y_{2}=(t_{1}^{1},t_{2}^{2}),\dots,y_{M}=(y_{1}^{1},t_{2}^{M}),y_{M+1}=(y_{1}^{2},y_{2}^{1}),\dots$ (2.11)
And we write $\mathcal{T}$ for the set of all profiles, $\mathcal{T}=\{y_{1},y_{2},\dots,y_{KM}\}$. Meanwhile, suppose that firm 1’s active type is 1, then the set of type profiles restricted to type $t_{1}^{1}$ is equal to
$\mathcal{T}_{1}^{1}=\{(t_{1}^{1},t_{2}^{1}),(t_{1}^{1},t_{2}^{2}),(t_{1}^{1},t_{2}^{3}),\dots,(t_{1}^{1},t_{2}^{M})\}$ (2.12)
@@
Definition 2.10 (Specific best-response strategy).
Suppose that firm $i$ ($i=1,2$) is of Type $k$. We say that $s_{i}^{k}$ is a best response strategy to the normalized strategy $s_{\sm i}\in\mathcal{S}_{\sm i}^{*}$ if:
$E_{i}^{k}(s_{i}^{k},s_{-i}^{*})\in\max_{s\in S_{i}^{k}}E_{i}^{k}(s,s_{-i}^{*})$ (2.15)
That is, $s_{i}^{k}$ is a best-response if it provides the highest possible payoff among all admissible strategies.
Definition 2.11 (Best-response strategy: normalized).
Let $s_{i}^{*}\in\mathcal{S}_{i}^{*}$ and $s_{-i}^{*}\in S_{-i}^{*}$ be normalized strategies. Then $s_{i}^{*}$ is a best-response to $s_{-i}^{*}$ if each component $s_{i}^{k}$ for $s_{i}^{*}$ is a best-response (see Definition 2.10 to $s_{-i}^{*}$.
@@
Two normalized strategies $s_{1}^{*},s_{2}^{*}$ for players 1 and 2 respectively are a Bayes equilibrium set of strategies if $s_{1}^{*}$ is a best-response to $s_{2}^{*}$ and $s_{2}^{*}$ is a simultaneously a best-response to $s_{1}^{*}$. In other words:
$s_{1}^{k}$ $\in\max_{s_{1}\in\mathcal{S}_{1}^{k}}E_{1}^{k}(s_{1},s_{2}^{*})$ (2.16a)
$s_{2}^{m}$ $\in\max_{s_{2}\in\mathcal{S}_{2}^{m}}E_{2}^{m}(s_{2},s_{1}^{*})$ (2.16b)
@@ 3.3 Market impact models
    Admissible strategies: Write $\mathcal{S}_{1}^{k}$ for the set of admissible strategies for firm 1, Type $k$, and similarly $\mathcal{S}_{2}^{m}$ for admissible strategies for firm 2, Type $m$;
@@
    Write $s_{1}^{k}\in\mathcal{S}_{1}^{k}$ and $s_{2}^{m}\in\mathcal{S}_{2}^{m}$ for admissible strategies for firms 1 and 2 showing up as types $k,m$ respectively;
@@
    Market impact parameter: Write $\kappa_{1}^{k},\kappa_{2}^{m}$ for the market impact parameters for firms 1 and 2 showing up as showing up as types $k,m$ respectively; and
@@
    Target quantities: Write $f_{1}^{k},f_{2}^{m}$ for the target quantities for firms 1 and 2 showing up as showing up as types $k,m$ respectively.
@@
    Strategy profiles: As in Definition 2.5 a normalized strategy for firms 1 is a strategy specification for each possible type of firm 1: $s_{1}^{*}=(s_{1}^{1},s_{2}^{2},\dots,s_{1}^{M})$. Similarly, a normalized strategy for firm 2 is a profile $s_{2}^{*}=(s_{2}^{1},\dots,s_{2}^{K})$,
@@ 3.5 Key assumptions
Every game has a set of valid pure strategies for each player. In trading games, the strategies are in the form of functions. The set of valid strategies are the admissible strategies. We assume that the sets of admissible strategies for firms 1 and 2 respectively consist of the set of twice-differentiable functions
@@
Assuming firm 1 is type $k$ and firm 2 is type $m$ then their conditional payoff functions are given by the negative values of the cost functions in Eq. 3.3.
@@
    The parameter $\kappa_{i}^{k}$ is referred to as the market impact parameter. The form of the conditional payoff functions means that each firm’s active type has a belief concerning what $\kappa$ is. These types are used by each active type to select their trading strategies. However, once they trade, both firms are subject to the same market conditions. There is only one market impact parameter governing trading costs. We call this the real market impact parameter. This will be discussed more in Section 4 when we discuss trading cost evaluation.

### v2

@@ 1 Introduction
In some circumstances, trading is a strategic interaction between market participants who are trading the same asset over the same period of time. This is the case where firms are concerned with the collective impact of their trades on market prices and choose how and when to trade in order to minimize their overall cost of trading, taking into consideration what the other firms are likely to do. The general situation just described, which we call trading in competition, was studied in [Chr24b] and [Chr24a] where it was shown that there unique Nash equilibria always exist (in the sense of [Nas50], [Nas51] and Definition 2.2 below) that only depend on the number of firms involved, their target quantities, and the market impact model the firms use to select their trades.
A key assumption and in our view a major shortcoming of our previous work is that it implicitly assumes that trading in competition is a game with complete information, one for which participants are in possession of full knowledge of all relevant aspects of the game, which in this context means the number of firms, their target quantities and the market impact model they use. By contrast, in most real-world situations, trading in competition is a game with incomplete information, in which some or all participants lack full knowledge of at least some aspects of the game. In the case of trading, firms do not know their competitors’ target quantities, and are far from sure of precise values of the market impact parametersor even which models they use.
@@
For more details on von Neumann and Morgenstern’s treatment of incomplete information, see [Mye04]. The main point of this discussion is that incomplete information is a persistent feature in trading situations due to institutional, regulatory and competitive factors.
The present paper aims to demonstrate how uncertainty can be incorporated into competitive trading using the framework of games with "Bayesian" players (or Bayesian games), as developed in John Harsanyi’s seminal work on games with incomplete information [Har67, Har94]. Harsanyi’s contributions provide a general framework for addressing information uncertainty in game theory and extend Nash’s equilibrium concept.
Building on this foundation, we extend the work on trading in competition presented in [Chr24a] and [Chr24b] by allowing for incomplete information. To achieve this, we first review the theory of games with Bayesian players in Section 2, where we detail the methods Harsanyi introduced to handle incomplete information. We then apply these concepts to the context of competitive trading.
In Section 3.1, we revisit the fundamental game-theoretic concepts employed in [Chr24a] and [Chr24b], providing the necessary groundwork for our analysis. Section 2 introduces the framework of Harsanyi’s Bayesian games, setting the stage for an in-depth exploration of incomplete information within the trading context. A central result of this approach is a method for transforming a game with incomplete information into an equivalent game of complete information. This transformation leverages Harsanyi’s framework to ensure that the resulting game retains the essential features of the original, but can be analyzed as a complete information game.
The advantage of this transformation is significant: complete information games are guaranteed to have at least one Nash equilibrium [^1] It is a fundamental result that all $n$-player non-cooperative complete information games have at least one Nash equilibrium [Nas51].. The equilibria derived in the transformed game can then be mapped back to the original incomplete information game, yielding strategies that are not only viable, but also exhibit properties interpretable as optimal within the context of competitive trading.
@@
The main result of this paper is a description of $n$-firm equilibria in trading in games with incomplete information, building on the framework introduced in [Chr24a]. This framework is based on the following assumptions, which are reviewed in detail in Section 3.3:
@@
    There are $n$ firms that trade in competition with each other over a fixed and common period of time (the "trading period").
@@
    Each firm uses the market impact model from the Almgren-Chriss framework (see [AC01]), a linear model with both permanent and temporary impact [^2] The model is linear in the sense that the impact on prices is linear in the quantity traded and therefore the cost of trading quadratic is quadratic in quantity traded.. This model features a single market impact parameter, referred to as the market impact coefficient.
The novelty of this paper lies in relaxing a key assumption implicit in [Chr24a], namely that each firm’s target quantity and market impact parameter are common knowledge. In game theory, a fact is considered common knowledge among players if (i) every player knows the fact, (ii) every player knows that every other player knows the fact, (iii) every player knows that every other player knows that every player knows the fact, and so on. The concept of common knowledge was explicitly introduced to game theory only in the 1970s [^3] The term ”common knowledge” may have originated in David Lewis’ philosophical treatise [Lew08] and was formally defined in game theory by Aumann in his seminal work [Aum76], though it can be argued that the idea was implicitly present in earlier works, such as those of von Neumann and Morgenstern..
@@ 2.3 Bayesian players and player types
Consider an $n$-player game where player $i$ ($i=1,\dots,n$) has one of $K_{i}$ types. Then we make the following definitions:
@@
    Type profile: A type profile is an $n$-tuple $y=\{t_{1}^{k_{1}},\dots,t_{n}^{k_{n}}\}$, associating to each player $i$ ($i=1,\dots,n$) a type $k_{i}$. Write $t_{i}^{k}\in y$ if player $k$ is type $i$ in the profile $y$.
@@
    Profile set: For a particular game, we write $\mathcal{T}$ for the set of all type profiles for the game.
@@
    Restricted profile set: Fix a player $i$ assume their type is $k$. Write $\mathcal{T}_{i}^{k}$ for the set of all type profiles $y_{i}^{k}\in\mathcal{T}$ such that $t_{i}^{k}\in y_{i}^{k}$.
@@ 2.4 Strategy profiles and payoff functions
We begin by defining the correct analogy to a strategy in a game with incomplete information. In a game with complete information a player faces a known adversary, which in this context means that a given player knows the type of their adversary. However, in a game with an incomplete information a player faces one of many possible types, each showing up with a given probability. Further, in a game with complete information, a player can determine the optimal action to take conditional on their adversary chooses to play a specific strategy. In a game with incomplete information, a player must choose the optimal strategy given their adversary would play a particular strategy when showing up as a particular type. We leverage this idea to define the notion of a strategy profile.
@@
    Strategy profile for a player: When player $i$ ($i=1,\dots,n$) is type $k$ then write $s_{i}^{k}$. Write
    $s_{i}^{*}=(s_{i}^{1},\dots,s_{i}^{K_{i}})$ (2.6)
    for the tuple of $K_{i}$ strategies thus represented. We call this a strategy profile for player $i$ and this represents a collection of conditional strategies, one for each type $k=1,\dots,K_{i}$ that player $i$ might show up as [^9] This was called a normalized strategy in [Har67]..
@@
    Strategy profile for all types in the game: Given a strategy profile $s_{i}^{*}$ for each player $i$ ($i=1,\dots,n$), define $s^{*}=(s_{1}^{*},\dots,s_{n}^{*})$ as the combined strategy profile for all players across all types.
@@
    Strategies: Each player is equipped with strategies $s_{i}^{*}$ consisting of a tuple $(s_{i}^{1},\dots,s_{i}^{K_{i}}$) such that $s_{i}^{k}$ is an admissible strategy for player $i$ when represented as type $k$, We assemble all of the strategy profiles into a single strategy profile $s^{*}$. Thus each player’s strategy in this game is a strategy profile $s_{i}^{*}$ consisting of admissible strategies in the original game;
@@ 2.5 Bayesian equilibrium for two firms
We now specialize type theory and Bayesian equilibrium to trading in competition and restrict ourselves to the case of two firm, illustrating how the theory works in the simpler case. Let $i=1,2$ denote the firms and let $K$ and $M$ be the number of admissible types for firms 1 and 2 respectively. For a given firm $i$, ${\sm i}$ will denote the "other firm", that is, if $i=1$, ${\sm i}=2$ and for $i=2$, ${\sm i}=1$. We write $\pi_{i}^{k}(m)$ for the probability that firm ${\sm i}$ is type $m$ conditional on firm $i$ being type $k$. Given this, the strategy profiles for firms as in definition 2.5 now simplify to a $K$-tuple of admissible strategies for firm 1 and an $M$-tuple for firm 2 as follows:
@@
and a strategy profile is simply the concatenation of $s^{*}$ of $s_{1}^{*}$ and $s_{2}^{*}$:
@@
Type profiles as in Definition 2.4 now simplify an assignment of a specific type to each firm. In simple terms, a type profile can be thought of as a specification of the actual type of each firm that "shows up" to trade. There are $K\times M$ type profiles, and these can be written in lexicographical order as follows:
$y_{1}=(t_{1}^{1},t_{2}^{1}),\,y_{2}=(t_{1}^{1},t_{2}^{2}),\,\dots,\,y_{M}=(t_{1}^{1},t_{2}^{M}),\,y_{M+1}=(t_{1}^{2},t_{2}^{1}),\dots,y_{K\cdot M}=(t_{1}^{K},t_{2}^{M})$ (2.12)
And we write $\mathcal{T}$ for the set of all profiles, $\mathcal{T}=\left\{y_{1},y_{2},\dots,y_{KM}\right\}$. Meanwhile, suppose that firm 1’s active type is 1, then the set of type profiles restricted to type $t_{1}^{1}$ is as follows:
$\mathcal{T}_{1}^{1}=\left\{(t_{1}^{1},t_{2}^{1}),\,(t_{1}^{1},t_{2}^{2}),\,(t_{1}^{1},t_{2}^{3}),\dots,\,(t_{1}^{1},t_{2}^{M})\right\}$ (2.13)
@@
Definition 2.10 (Specific best-response to a strategy profile).
When studying trading as a game of complete information, the notion of a best-response strategy is clear and intuitive. If, for example, firm $2$ trades the strategy $s_{2}$ then player 1’s best-response strategy to $s_{2}$ is the strategy that minimizes the expected cost when trading in competition with $s_{2}$ [^10] This is discussed in detail in [Chr24b]. In a game with incomplete information firm $i$ ($i=1,2$) faces one of many possible adversaries depending on which of the possible active types for firm $\sm i$ shows up.
In this case, to define a best response strategy firm 1 (say), we have to contend with the uncertainty as to which firm 2 active type shows up. Firm 1 faces not a strategy but a strategy profile $s_{2}^{*}=(s_{2}^{1},\dots,s_{2}^{M})$ (see Definition 2.5). Recalling that $S_{i}^{k}$ is the set of admissible strategies for firm $i$, type $k$. We say the specific strategy $s_{1}^{k}$ is the best-response to the strategy profile $s_{2}^{*}$ if its expected value $E_{1}^{k}(s_{1}^{k},s_{2}^{*})$, Eq. 2.14, is maximal among all strategies $s\in S_{1}^{k}$:
$E_{1}^{k}(s_{1}^{k},s_{2}^{*})=\max_{s\in S_{1}^{k}}E_{1}^{k}(s,s_{2}^{*})$ (2.16)
Likewise $s_{2}^{k}$ is a best-response to $s_{1}^{*}$ if
$E_{2}^{k}(s_{2}^{k},s_{2}^{*})=\max_{s\in S_{2}^{k}}E_{2}^{k}(s,s_{1}^{*})$ (2.17)
Definition 2.11 (Best-response strategy profile).
Now consider firm 1 from firm 2’s perspective. Suppose that firm 2 believes that firm 1 is trading the strategy profile $s_{1}^{*}$ and wants to produce a strategy profile $s_{2}^{*}$ each component of which is a best-response to $s_{1}^{*}$. Such a strategy profile is called a best-response strategy profile. A similar definition applies for the best response profile for firm 1 versus a known profile $s_{2}^{*}$ for firm 2. What is the meaning of this?
In that firm 2 (say) knows what type it is, why does it need a complete best-response profile? The answer has to do with firm 1’s perspective. Since firm 1 does not know what firm 2’s active type will be, it is necessary for firm 2 to have a best-response strategy for each of its potential types.
@@
Strategy profiles $s_{1}^{*},s_{2}^{*}$ for firms 1 and 2 respectively are a Bayes equilibrium set of strategies if $s_{1}^{*}$ is a best-response profile to $s_{2}^{*}$ and $s_{2}^{*}$ is simultaneously a best-response to $s_{1}^{*}$. In other words:
$s_{1}^{k}$ $=\max_{s_{1}\in\mathcal{S}_{1}^{k}}E_{1}^{k}(s_{1},s_{2}^{*}),\qquad k=1,\dots,K$ (2.18a)
$s_{2}^{m}$ $=\max_{s_{2}\in\mathcal{S}_{2}^{m}}E_{2}^{m}(s_{2},s_{1}^{*}),\qquad m=1,\dots,M$ (2.18b)
@@ 3.3 Market impact models
    Admissible strategies: Write $\mathcal{S}_{1}^{k}$ for the set of admissible strategies available to firm 1 when its active type is $k$. Similarly write $\mathcal{S}_{2}^{m}$ for admissible strategies for firm 2 when its active type is $m$;
@@
    Specific admissible strategies: Write $s_{1}^{k}\in\mathcal{S}_{1}^{k}$ and $s_{2}^{m}\in\mathcal{S}_{2}^{m}$ for specific admissible strategies for firms 1 and 2 when their active types are $k,m$ respectively;
@@
    Market impact coefficient: Write $\kappa_{1}^{k},\kappa_{2}^{m}$ for the market impact parameters used by firms 1 and 2 (respectively) when their active types are $k,m$ respectively; and
@@
    Target quantities: Write $f_{1}^{k},f_{2}^{m}$ for the target quantities for firms 1 and 2 when their active types are $k,m$ respectively.
@@
    Strategy profiles: As in Definition 2.5 a strategy profile for firm 1 is a strategy specification for each of its possible active types. This is written as an $M$-tuple of strategies $s_{1}^{*}=(s_{1}^{1},s_{1}^{2},\dots,s_{1}^{M})$. Similarly, a strategy profile for firm 2 is a a $K$-tuple of strategies $s_{2}^{*}=(s_{2}^{1},\dots,s_{2}^{K})$,
@@ 3.5 Key assumptions
We begin by assuming that every game has specified set of pure strategies available to each player. In trading games, the strategies are trading strategies specified by functions of time [^12] The details of such strategies are given in [Chr24b] and [Chr24a]. The set of valid strategies are the admissible trading strategies. We assume that the sets of admissible strategies for firms 1 and 2 respectively consist of the set of twice-differentiable functions
@@
Assume that firm 1’s active type is $k$ and firm 2’s is $m$ and that firm 1 trades the strategy $s_{1}^{k}$ when its active type is $k$, and firm 2 trades $s_{2}^{m}$ when its active type is $m$. Then when firm 1 competes with the strategy $s_{2}$ and firm 2 competes with the strategy $s_{1}$. we write $V_{1}^{k}(s_{1}^{k},s_{2}$ and $V_{2}^{m}(s_{1},s_{2}^{m})$ for the conditional payoff functions for firms 1 and 2 respectively. These are given by the negative values of the cost functions in Eq. 3.3:
@@
    The parameter $\kappa_{i}^{k}$ is referred to as the market impact coefficient. The form of the conditional payoff functions means that each firm’s active type has a belief concerning what $\kappa$ is. These types are used by each active type to select their trading strategies. However, once they trade, both firms are subject to the same market conditions. There is only one market impact parameter governing trading costs. We call this the real market impact parameter. This will be discussed more in Section 4 when we discuss trading cost evaluation.